OpenStudy (hocuspucus292333393):

What equation represents this description? When 3/4 is added to x, the result is 9/8 . Enter your answers in the boxes in simplest form.

OpenStudy (sunnnystrong):

x+3/4 = 9/8

OpenStudy (sunnnystrong):

x=1/4

OpenStudy (koolkid98):

i beleive she is right ^ |

OpenStudy (hocuspucus292333393):

it is right how i put it koolkid

OpenStudy (koolkid98):

like that